Lifetime savings accounts, known as LSAs, allow people to invest after-tax money without being taxed on any of the gains. IfHow much money would be accumulated 13 years from now from deposits of $10,000 per year for 5 consecutive years, starting 6 yHi, Could you solve/fill the blank? Thanks a lot
Answer #1

rate positively.

Ans 1 We have to use finanical calculator to solve this
put in calculator - (Set the calculator in BEGIN mode)
PV 0
PMT -16000
I 10%
N 15
Compute FV $508,359.71
ans = $508,359.71
Ans 2 Computation of future value
Year Investment Future value
6 10000     19,487.17 =10000*(1+10%)^7
7 10000     17,715.61 =10000*(1+10%)^6
8 10000     16,105.10 =10000*(1+10%)^5
9 10000     14,641.00 =10000*(1+10%)^4
10 10000     13,310.00 =10000*(1+10%)^3
Total     81,258.88
Answer =     81,258.88